OpenStudy (anonymous):

2x^2-4x=1 Pleasse help!

jimthompson5910 (jim_thompson5910):

do you know the quadratic formula?

OpenStudy (anonymous):

X=-,+ square root of 3/2+1

hartnn (hartnn):

write that as 2x^2-4x-1=0 Compare this quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)